# Towards Open-World Human Action Segmentation Using Graph Convolutional Networks

### v1

@@ III-C Temporal Clustering Loss
To enforce temporally consistent and discriminative feature clusters for open-world generalization, we propose the Temporal Clustering Loss, a distribution-aware contrastive objective inspired by supervised contrastive learning [23]. Unlike conventional contrastive losses that operate on individual samples, our formulation explicitly models class-wise distributions in the spatio-temporal feature space. This is critical for action segmentation, where intra-class temporal variability and inter-class similarity are key challenges.
Let $\mathcal{F}_{i}=\{f_{t}\}_{t=1}^{T}$ denote the temporal sequence of embeddings for class $i\in N$, where $f_{t}\in\mathbb{R}^{d}$ is there frame-level feature at time $t$.
@@
where $\mathcal{L}_{CE}$ is the cross-entropy loss, ensuring classification accuracy, $\beta$ balances the contributions. By jointly optimizing discriminative classification and geometrically structured embeddings, the model learns temporally stable features that generalize to unseen actions while avoiding overconfidence on outliers.
@@ III-D Mixup
We adopt the Mixup [24] data augmentation technique to enhance model robustness and generalization to Out-of-Distribution (OOD) samples by enforcing linear feature transitions between classes. Unlike traditional empirical risk minimization (ERM), which learns only from observed training samples, Mixup trains the model on convex interpolations of input-label pairs, explicitly regularizing the feature space geometry. For spatio-temporal skeleton sequences, this is critical as OOD actions often manifest as semantic interpolations between known classes (e.g., a mix of running and jumping).
@@
where $\lambda\sim\text{Beta}(\alpha,\alpha)$ and $\alpha=0.2$ controls the interpolation strength. Lower $\alpha$ skews $\lambda$ towards extremes (0 or 1), preserving semantic coherence in skeleton sequences while still expanding vicinal distributions. By leveraging Vicinal Risk Minimization (VRM), Mixup expands the learned feature space, mitigating overfitting and improving generalization compared to traditional Empirical Risk Minimization (ERM).
@@
$ACC_{1}$ $ACC_{2}$ $F1@10$ $F1@25$ $F1@50$ $AUROC$ $ACC_{\text{OOD}}$ $h_{\text{score}}$
PGCN [1] (Baseline) 79.61 70.39 83.70 81.88 73.10 50.00 $\mathbf{85.86}$ 63.21
@@
    1All configurations are evaluated on the Bimacs [6] subject 1 testset. The best results across all modifications are highlighted in bold.
@@ IV-A Dataset
The Bimanual Actions (Bimacs) [6] dataset comprises 540 RGB-D recordings (2 hours and 18 minutes) at 15 FPS, capturing bimanual tasks like pouring milk while stirring cereal in a kitchen and workshop. It includes 26 nodes (12 for human joints and 14 for object centers), and framewise annotations for 12 objects and 6 subjects, covering 14 action categories.
The 2 Hands and Object (H2O) [7] dataset contains 571,645 frames from 4 participants performing 37 actions in environments like a hall, office, and kitchen. It includes 42 hand pose nodes and 8 object joints, capturing actions like grabbing, placing, and pouring. The dataset is recorded at 30 FPS with synchronized RGB and depth images using multiple cameras mounted on a headset.
@@ IV-B Experimental settings
TABLE II: Comparison of the frame-wise performance and open-set F1@K score of our method against other state-of-the-art frameworks. Bimacs rows correspond to results on the Bimanual Actions dataset [6], while H2O rows correspond to the 2 Hands and Object dataset [7].
Dataset Methods Closed-Set Open-Set Out-of-Distribution
ACCclose ACCopen $F1@10$ $F1@25$ $F1@50$ $AUROC$ $ACC_{OOD}$ $h_{score}$
Bimacs [6] PGCN [1] 79.61 70.39 83.70 81.88 73.10 50.00 $\mathbf{85.86}$ 63.21
ST-GCN+TPP [2] 74.16 65.77 88.62 86.73 77.35 50.00 35.26 41.36
AGCN+TPP [11] 73.93 65.62 85.63 83.21 72.06 51.40 36.65 42.79
CTR-GCN+TPP [25] 74.11 65.44 89.97 88.49 65.44 50.00 44.62 47.16
UQ-TFGCN [5] 83.38 74.98 90.25 88.28 78.83 51.79 69.77 59.25
UQ-TFGCN + mixup + $\mathcal{L}_{TC}$ 84.44 81.94 93.72 92.76 87.10 69.78 69.78 69.78
EPGCN $\mathbf{85.21}$ $\mathbf{85.71}$ $\mathbf{95.10}$ $\mathbf{95.08}$ $\mathbf{90.03}$ $\mathbf{84.62}$ 84.69 $\mathbf{84.65}$
H2O [7] PGCN [1] 81.72 61.17 85.58 80.29 $\mathbf{74.02}$ 50.00 47.75 48.85
ST-GCN+TPP [2] 72.99 54.08 76.24 70.49 55.28 50.00 52.37 51.57
AGCN+TPP [11] 79.48 61.95 83.97 78.52 70.45 50.00 61.38 55.11
UQ-TFGCN [5] $\mathbf{88.09}$ 65.51 $\mathbf{88.56}$ $\mathbf{82.69}$ 72.83 50.00 66.10 56.93
EPGCN 81.95 $\mathbf{74.94}$ 86.50 80.00 73.06 $\mathbf{72.97}$ $\mathbf{84.28}$ $\mathbf{78.22}$
@@ IV-C Ablation studies
We systematically evaluate the impact of design choices on open-world performance using the Bimanual Actions dataset [6]. Starting with the PGCN baseline [1], we incrementally introduce components to isolate their contributions.
@@ IV-D Comparison with the state-of-the-art
The proposed EPGCN framework is compared with state-of-the-art action segmentation framework on the Bimacs [6] and H2O [7] datasets. Several popular graph convolutional networks: ST-GCN [2], AGCN [11], CTR-GCN [25], and UQ-TFGCN [5]. ST-GCN, AGCN and CTR-GCN are combined with the temporal pyramid pooling (TPP) decoder module since they are not originally designed for fine-grained segmentation task.
Table II presents the results on all three tasks: close-set recognition, open-set segmentation, and Out-of-Distribution classification, where the top and bottom halves correspond to the performance on the Bimacs and H2O dataset, respectively. Its seen that our EPGCN framework outperforms all other frameworks on both datasets. UQ-TFGCN, which incorporates spectral normalization in its residual layers to preserve feature-space distances for OOD detection, ranks second. Notably, EPGCN improves upon UQ-TFGCN in F1@50 and $h_{\text{score}}$ by 11.2% and 25.4%, respectively. This significant improvement stems from UQ-TFGCN’s focus on covariate shift, as it was evaluated on a noisy Bimacs dataset and the IKEA Assembly dataset [26], which is semantically dissimilar to Bimacs due to its 2D spatial representation rather than 3D. However, when UQ-TFGCN is trained with Mixup and our proposed Temporal Clustering Loss, F1@50 and $h_{\text{score}}$ increase by 8.3% and 10.5% compared to training with UQ-TFGCN alone. This confirms the effectiveness of our proposed components in improving generalization in open-world scenarios.
@@ V CONCLUSIONS
In this work, we introduce the Open-World Action Segmentation problem and propose a novel framework that addresses key limitations in existing open-world recognition methods. Unlike prior approaches that rely on incremental learning or external labeling, our method uses a distance-based classifier (K-means), assuming that Out-of-Distribution samples form distinct clusters without requiring manual labeling. We enhance the closed-set PGCN model by integrating a Temporal Efficient Upsampling (TEU) module, which better fuses encoder features across temporal dimensions, and apply Mixup to introduce uncertainty during training, helping the model generalize to various OOD scenarios. Additionally, Temporal Clustering Loss enhances the model’s ability to form more distinct clusters in the feature space. Evaluations on the Bimacs and H2O datasets show that our framework consistently outperforms existing methods (PGCN and UQ-TFGCN), demonstrating the effectiveness of these novel components for open-world action segmentation.
The OWAS problem aims to provide a streamlined approach for labeling unknown samples using only signals from known samples, without relying on external information, which can be challenging to obtain especially for dense pixel-wise labeling on new videos. This capability is particularly valuable in real-world robotics applications, where adapting to unseen actions without manual annotations is crucial for safe, reliable, and cost-effective operation.

### v2

@@ III-C Temporal Clustering Loss
To enforce temporally consistent and discriminative feature clusters for open-world generalization, we propose the Temporal Clustering Loss, a distribution-aware contrastive objective inspired by supervised contrastive learning [khosla2020supervised]. Unlike conventional contrastive losses that operate on individual samples, our formulation explicitly models class-wise distributions in the spatio-temporal feature space. This is critical for action segmentation, where intra-class temporal variability and inter-class similarity are key challenges.
Let $\mathcal{F}_{i}=\{f_{t}\}_{t=1}^{T}$ denote the temporal sequence of embeddings for class $i\in N$, where $f_{t}\in\mathbb{R}^{d}$ is the frame-wise feature at time $t$.
@@
where $\mathcal{L}_{CE}$ is the cross-entropy loss, ensuring classification accuracy, $\beta$ balances the contributions. By jointly optimizing discriminative classification and geometrically structured embeddings, the model learns temporally stable features that generalize to novel actions while avoiding overconfidence on outliers.
@@ III-D Mixup
We adopt the Mixup [Zhang2017mixupBE] data augmentation technique to enhance model robustness and generalization to Out-of-Distribution (OOD) samples by enforcing linear feature transitions between classes. Unlike traditional empirical risk minimization, which learns only from observed training samples, Mixup trains the model on convex interpolations of input-label pairs, explicitly regularizing the feature space geometry. For spatio-temporal skeleton sequences, this is critical as OOD actions often manifest as semantic interpolations between known classes (e.g., a mix of running and jumping).
@@
where $\lambda\sim\text{Beta}(\alpha,\alpha)$ and $\alpha=0.2$ controls the interpolation strength. Lower $\alpha$ skews $\lambda$ towards extremes (0 or 1), preserving semantic coherence in skeleton sequences while still expanding vicinal distributions. By leveraging Vicinal Risk Minimization, Mixup expands the learned feature space, mitigating overfitting and improving generalization compared to traditional Empirical Risk Minimization.
@@
$ACC_{1}$ $ACC_{2}$ $F1@10$ $F1@25$ $F1@50$ AUROCAUROC $ACC_{\text{OOD}}$ $h_{\text{score}}$
PGCN [xing2022understanding] (Baseline) 79.61 70.39 83.70 81.88 73.10 50.00 $\mathbf{85.86}$ 63.21
@@
    1All configurations are evaluated on the Bimacs [Dreher2019LearningOR] subject 1 testset. The best results across all modifications are highlighted in bold.
@@ IV-A Dataset
The Bimanual Actions (Bimacs) [Dreher2019LearningOR] dataset comprises 540 RGB-D recordings (2 hours and 18 minutes) at 15 FPS, capturing bimanual tasks like pouring milk while stirring cereal in a kitchen and workshop. It includes 26 nodes (12 for human joints and 14 for object centers), and framewise annotations for 12 objects and 6 subjects, covering 14 action categories.
The Two Hands and Object (H2O) [Kwon_2021_ICCV] dataset contains 571,645 frames from 4 participants performing 37 actions in environments like a hall, office, and kitchen. It includes 42 hand pose nodes and 8 object joints, capturing actions like grabbing, placing, and pouring. The dataset is recorded at 30 FPS with synchronized RGB and depth images using multiple cameras mounted on a headset.
@@ IV-B Experimental settings
TABLE II: Comparison of the frame-wise performance and open-set F1@K score of our method against other state-of-the-art frameworks. Bimacs rows correspond to results on the Bimanual Actions dataset [Dreher2019LearningOR], while H2O rows correspond to the 2 Hands and Object dataset [Kwon_2021_ICCV].
Dataset Methods #\#Paramb Closed-Set Open-Set Out-of-Distribution
ACCclose ACCopen $F1@10$ $F1@25$ $F1@50$ AUROCAUROC $ACC_{OOD}$ $h_{score}$
Bimacs PGCN [xing2022understanding] 5.42 M 79.61 70.39 83.70 81.88 73.10 50.00 $\mathbf{85.86}$ 63.21
ST-GCN+TPP [Yan2018SpatialTG] 5.05 M 74.16 65.77 88.62 86.73 77.35 50.00 35.26 41.36
AGCN+TPP [2sagcn2019cvpr] 5.42 M 73.93 65.62 85.63 83.21 72.06 51.40 36.65 42.79
CTR-GCN+TPP [9710007] 3.36 M 74.11 65.44 89.97 88.49 65.44 50.00 44.62 47.16
UQ-TFGCN [xing2024understanding] 20.13 M 83.38 74.98 90.25 88.28 78.83 51.79 69.77 59.25
UQ-TFGCN+b 20.13 M 84.44 81.94 93.72 92.76 87.10 69.78 69.78 69.78
EPGCN 5.64 M $\mathbf{85.21}$ $\mathbf{85.71}$ $\mathbf{95.10}$ $\mathbf{95.08}$ $\mathbf{90.03}$ $\mathbf{84.62}$ 84.69 $\mathbf{84.65}$
H2O PGCN [xing2022understanding] 5.42 M 81.72 61.17 85.58 80.29 $\mathbf{74.02}$ 50.00 47.75 48.85
ST-GCN+TPP [Yan2018SpatialTG] 5.05 M 72.99 54.08 76.24 70.49 55.28 50.00 52.37 51.57
AGCN+TPP [2sagcn2019cvpr] 5.42 M 79.48 61.95 83.97 78.52 70.45 50.00 61.38 55.11
UQ-TFGCN [xing2024understanding] 20.13 M $\mathbf{88.09}$ 65.51 $\mathbf{88.56}$ $\mathbf{82.69}$ 72.83 50.00 66.10 56.93
EPGCN 5.64 M 81.95 $\mathbf{74.94}$ 86.50 80.00 73.06 $\mathbf{72.97}$ $\mathbf{84.28}$ $\mathbf{78.22}$
@@
  • b
    b The number of learnable parameters.
  • c
    c The original UQ-TFGCN model is trained with the mixup data augmentation and the $\mathcal{L}_{TC}$ loss.
@@ IV-C Ablation studies
We systematically evaluate the impact of design choices on open-world performance using the Bimanual Actions dataset [Dreher2019LearningOR]. Starting with the Pyramid Graph Convolutional Network (PGCN) baseline [xing2022understanding], we incrementally introduce components to isolate their contributions.
@@ IV-D Comparison with the state-of-the-art
The proposed EPGCN framework is compared with state-of-the-art action segmentation framework on the Bimacs [Dreher2019LearningOR] and H2O [Kwon_2021_ICCV] datasets. Several popular graph convolutional networks: ST-GCN [Yan2018SpatialTG], AGCN [2sagcn2019cvpr], CTR-GCN [9710007], and UQ-TFGCN [xing2024understanding]. ST-GCN, AGCN and CTR-GCN are combined with the temporal pyramid pooling (TPP) decoder module since they are not originally designed for fine-grained segmentation task.
Table II presents the results on all three tasks: close-set recognition, open-set segmentation, and Out-of-Distribution classification, where the top and bottom halves correspond to the performance on the Bimacs and H2O dataset, respectively. Its seen that our EPGCN framework outperforms all other frameworks on both datasets. UQ-TFGCN, which incorporates spectral normalization in its residual layers to preserve feature-space distances for OOD detection, ranks second. Notably, EPGCN improves upon UQ-TFGCN in F1@50 and $h_{\text{score}}$ by 11.2% and 25.4%, respectively. This significant improvement stems from UQ-TFGCN’s focus on covariate shift, as it was evaluated on a noisy Bimacs dataset and the IKEA Assembly dataset [BenShabat2020TheIA], which is semantically dissimilar to Bimacs due to its 2D spatial representation rather than 3D. However, when UQ-TFGCN is trained with Mixup and our proposed Temporal Clustering Loss, F1@50 and $h_{\text{score}}$ increase by 8.3% and 10.5% compared to training with UQ-TFGCN alone. This confirms the effectiveness of our proposed components in improving generalization in open-world scenarios.
@@ V CONCLUSIONS
In this work, we formalize the Open-World Action Segmentation (OWAS) problem and propose a novel framework that addresses critical limitations in existing open-world recognition. Our approach employs a distance-based classifier (K-means) to automatically cluster novel actions without manual labeling, leveraging the observation that Out-of-Distribution (OOD) samples form distinct groupings. We enhance a closed-set backbone through three key innovations: (1) A Temporal Efficient Upsampling (TEU) module that fuses multi-scale encoder features for improved temporal coherence, (2) Mixup-based uncertainty injection during training to generalize across diverse OOD scenarios, and (3) Temporal Clustering Loss to enforce separable feature representations. Evaluations on public datasets demonstrate state-of-the-art performance, validating our framework’s efficacy for segmenting both known and novel actions.
While this work provides a streamlined solution for labeling novel actions using only signals from known classes, it is limited by the clustering algorithm K-means, which requires predefined number of classes. Future work will focus on continual cluster refinement for seamless novel action integration, cross-modal generalization through motion data fusion to enhance real-world robustness, risk-aware novelty detection to prevent high-stakes misclassification, and further deploy the framework in safety-critical Human-Robot Interaction scenarios.
ACKNOWLEDGMENT
The research has been (partially) supported by the German Research Foundation (DFG), SFB 1320 EASE, CRC, University of Bremen. The research was conducted in subproject R1: NEEM-based embodied knowledge system.